OpenStudy (anonymous):

A company has a machine that can complete a job in 6 hours. With the help of a newer machine, the job can be completed in 2 hours. How long would the new machine take to do the job alone?

jimthompson5910 (jim_thompson5910):

\[\large \frac{1}{6}+\frac{1}{x}=\frac{1}{2}\] \[\large \frac{x}{6x}+\frac{1}{x}=\frac{1}{2}\] \[\large \frac{x}{6x}+\frac{6}{6x}=\frac{1}{2}\] \[\large \frac{x+6}{6x}=\frac{1}{2}\] \[\large 2(x+6)=6x\] \[\large 2x+12=6x\] \[\large 12=6x-2x\] \[\large 12=4x\] \[\large \frac{12}{4}=x\] \[\large 3=x\] \[\large x=3\] So it would take 3 hours if the new machine was working alone.